OpenStudy (anonymous):

\[\int\limits \sqrt(x) / (1+ x^.25) \]

OpenStudy (anonymous):

\[\int\sqrt{\frac{x}{1+x^{1/4}}}~dx~~?\] or \[\int\frac{\sqrt x}{1+x^{1/4}}~dx~~?\]

OpenStudy (anonymous):

the 2nd one!

OpenStudy (anonymous):

Notice that \(\left(x^{1/4}\right)^2=x^{2/4}=x^{1/2}=\sqrt x\). So if you let \(u=x^{1/4}\), so that \(u^2=x^{1/2}\). Differentiating, you get \(du=\dfrac{1}{4}x^{-3/4}~dx\), or \(4x^{3/4}~du=dx\), or \(4u^3~du=dx\). So the integral becomes \[\int\frac{u^2}{1+u}~(4u^3~du)\\ 4\int\frac{u^5}{1+u}~du\]

OpenStudy (anonymous):

Thank you, I see now where I messed up the problem! Appreciate your help

OpenStudy (anonymous):

You're welcome!